# Friedrich-Alexander-Universität Erlangen-Nürnberg, Erlangen, Germany

### v1

@@ Abstract
Analyzing the cone photoreceptor pattern in images obtained from the living human retina using quantitative methods can be crucial for the early detection and management of various eye conditions. Confocal adaptive optics scanning light ophthalmoscope (AOSLO) imaging enables visualization of the cones from reflections of waveguiding cone photoreceptors. While there have been significant improvements in automated algorithms for segmenting cones in confocal AOSLO images, the process of labelling data remains labor-intensive and manual. This paper introduces a method based on deep learning (DL) for detecting and segmenting cones in AOSLO images. The models were trained on a semi-automatically labelled dataset of 20 AOSLO batches of images of 18 participants for 0, 1, and 2 from the foveal center. F1 scores were 0.968, 0.958, and 0.954 for 0, 1, and 2, respectively, which is better than previously reported DL approaches. Our method minimizes the need for labelled data by only necessitating a fraction of labelled cones, which is especially beneficial in the field of ophthalmology, where labelled data can often be limited.
@@ 1 Introduction
Considering just the cones, peak foveal density can approach 200,000 cones per $mm^{2}$ [5], making manual labelling impractical. On the other hand, existing automatic labelling techniques may not consistently enable the automatic identification of every cone within an image, particularly in the presence of blood vessels or when the image clarity is compromised. Furthermore, the challenge intensifies when examining retinal locations that are more eccentric from the fovea.
@@ 3.1 Dataset
In this study, we employed a semi-automatically labelled dataset of 20 AOSLO batches of images of 18 participants with normal vision, who belong to a wide variety of age and health groups and represent a wide range of retinas. Each batch consists of approximately 40 confocal images. The cone centers were first automatically identified with the classical method [7]. Then, missed cones were manually added by a human expert or removed if they were mislabelled by the automatic algorithm, representing approximately 5% of all cones across the dataset images. The data was split on the participant’s level so that images belonging to one participant appear only in one subset, with the train:test split ratio of 70:30. Therefore, we had 14 batches (540 images) of AOSLO images for training and 6 batches (240 images) for testing. Each image was cropped to $550\times 550$ pixel resolution with 350 labelled cones on each image on average, for a total of 190k segmented cells in the training subset at the starting point. Fig. 1 shows examples of confocal images with labelled cone centers.
Figure 1: Two examples of confocal AOSLO images with labelled cone centers using the existed semi-automatic segmentation method [7] with the following refinement by a medical expert.
@@ 3.2 Human-in-the-loop approach
Fig. 2 illustrates the overall pipeline of the proposed method. AOSLO images were labelled and split into the test and train subsets as described in the dataset section 3.1. To all labelled areas, we applied the Voronoi algorithm to obtain the masks of the cones. Then, the human-in-the-loop step was applied: on the initially labelled AOSLO images, we trained DL-based models to generate semantic masks on unlabelled images. Then, from semantic masks, we calculated the center of mass for each segment (cone), which is basically the center of the cone we manually labelled. Therefore, we could evaluate the models by comparing the obtained centers with ground truth labels. Given the potential for initial inaccuracies, manual correction is a crucial step in the method. This step ensures the precision of the model’s output by allowing for the expert review and adjustment of the segmented cone centers, mitigating the risk of errors in the initial automated segmentation. Adding new annotations and the manual correction step are not involved in the initial zero iteration, they are only applied starting from the first iteration step.
The Voronoi algorithm is reapplied to the refined data after correction of the centers of the cones, which was done in EXACT [9]. This iterative process refines the segmentation accuracy and enriches the training dataset with additional, corrected instances. Thus, the next 15% of the total number of images is labelled at each iteration, increasing the training dataset. Each iteration concludes with an evaluation step on a test dataset to quantify the improvements. Additionally, at this stage, we apply the $K$-means algorithm for clustering the cones by the mean brightness of the center part (reflection). Therefore, we obtain the percentage of reflecting and non-reflecting cones, which is also a priori information for diagnostics. This cyclical process, encompassing both automated segmentation and expert review, ensures the development of a robust model capable of high-precision cone segmentation.
Figure 2: Pipeline of the method. Voronoi algorithm is applied to initially semi-automatically annotated cones to obtain the masks. Then, a segmentation model was trained, which generates segmentation masks for unlabelled AOSLO images. The center of mass function is applied to get the centers of the cells from segmentation masks. After each iteration step, the model was evaluated using the test subset. A manual correction step is involved in the pipeline to improve the annotations of the segmentation model of initially unlabelled images.
@@ 3.3 Voronoi algorithm
The Voronoi algorithm is one of the more useful geometrical constructions to study point patterns since it provides all the information needed to study proximity relations between points [10]. Connecting surrounding cones and characterizing the number of sides, the Voronoi diagram allows assessment of the degree of hexagonality, and it is often used to show how disease and aging can affect this aspect of packing geometry [1]. In a healthy retina, cones are packed in the most efficient manner possible, which is a hexagonal (honeycomb) arrangement. The degree of hexagonality, therefore, can be used as a proxy for general retinal health. Applying the Voronoi algorithm, we can obtain a reasonably accurate approximation of photoreceptor segmentation by labelling only the centers of cones.
Figure 3: Application of Voronoi algorithm on the labelled AOSLO images: (a) example of original image; (b) segmented image.
@@ 3.4 Attention-augmented U-Net
Fig. 4 shows the overview of the model. In our model, we applied the concept of flows (vector gradient fields) [20]. This means that we trained a neural network to predict the horizontal and vertical gradients of the topological maps. Additionally, the network predicts a binary map to indicate if a given pixel is inside or outside of regions of interest. Our model was based on the general U-Net architecture [14] with an additional attention-augmented module (AA module) [13]. This module dynamically adjusts the importance of different spatial regions and channels in the input data, enabling the network to prioritize more relevant features for improved segmentation accuracy. Such augmentation facilitates precise localization and detailed segmentation in complex image datasets, which is particularly beneficial in medical imaging applications where accuracy is essential. AOSLO images have noise distributed evenly across the image. Attention mechanisms can help the model to focus on relevant features and ignore distractions, therefore, improving segmentation accuracy.
@@
Figure 4: Model overview. (a) Transformation from the center of the cell to a gradient vector field using Voronoi algorithm. (b) U-Net model with additional Attention-augmented module.
@@ 4 Experiments
The model was trained for 500 epochs on each iteration with stochastic gradient descent with a learning rate of 0.001, a momentum of 0.9, a batch size of 16 images, and a weight decay of 0.0001. All the models were trained on a single NVIDIA A100 graphics processing unit on a machine with two Intel Xeon Gold 6134 3.2 GHz and 96 GB RAM.
To predict the horizontal and vertical gradients, we used the L2 loss function. We applied the cross-entropy loss function to predict the probability that a pixel was inside or outside a cell.
@@ 4.2 Metrics
where $a$ and $b$ are predicted and ground truth center, respectively. Detected cones were evaluated using Recall, Precision, and F1-score:
@@ 4.3 Results
Fig. 5 shows an example of the application of the algorithm on the first (a) and second (b) iterations. Green and red dots correspond to ground truth and predicted cell centers, respectively. Blue squares on the first iteration show unlabelled cells (FNs); on the second iteration, they are correctly labelled.
@@
Figure 5: Evaluation of the proposed method on first (a) and second (b) iterations on a 0 test sample. Green dots correspond to the ground truth cone centers, and red to the predicted centers. Blue squares show the False Negative predictions of the model.
@@
Table 1 presents the comparative performance of the StarDist, Cellpose, and our models. Recall, Precision, and F1 score were computed separately for 0, 1, and 2 from the fovea. $D_{L_{2}}$ was calculated for all three degrees together. All models are characterized by a fundamental improvement after the second correction: on average, the F1 score improved by 7%. We also see a tendency for the scores to deteriorate slightly as one moves away from the fovea.
The obtained centers were clustered in terms of brightness to monitor the distribution of light-reflecting and dark photoreceptors during training. The $K$-means algorithm is a popular unsupervised machine learning technique for clustering data into a specified number of clusters, denoted by $K$. We applied the $K$-means algorithm during each iteration for additional distribution of reflecting cone control. Clustering the cones based on their brightness level is particularly useful in retinal imaging for understanding differences between healthy and diseased retinas. Fig. 7 shows examples of clustering on images of 0, 1, and 2.
@@
Figure 7: Examples of the method performance on the test images located at 0, 1, and 2 from the fovea with applied $K$-Means clustering algorithm. Blue-marked labels correspond to the cones with the highest reflection, green to the lowest, and red to the mid. White boxes show the location of the zoomed area on the right column.
@@
Figure 8: The cumulative average number of corrections of cone centers on the first and second iterations per image for Cellpose, StarDist, and the proposed model. The initial training iteration was done with the original labelled dataset; therefore, the number of corrections was equal to zero.
@@ 5 Conclusion
This work describes and evaluates a method for the identification and segmentation of cone photoreceptors from AOSLO confocal images. Models were trained and tested on images covering a more extensive range of images of 18 participants with only 5% labelled cones. Our proposed method received an overall F1 score of 0.968 for cones for 0, 0.958 for 1, and 0.954 for 2, which is better than previously reported DL approaches [4, 3]. Our method can reduce the labelling effort by requiring only a portion of labelled cones and is particularly advantageous in the ophthalmology field, where labelled data can be scarce. The work is limited to the range of degrees of the distance from the center from the fovea — 0, 1, 2. Rods are already present at 1 but peaks in density at around 15, thus rods become more and more visible on images between cones, which also require detection. Therefore, a potential improvement of the method could be to add the annotations for rods and classify cones and rods for eccentricities more than 2.

### v2

@@ Abstract
Analyzing the cone photoreceptor pattern in images obtained from the living human retina using quantitative methods can be crucial for the early detection and management of various eye conditions. Confocal adaptive optics scanning light ophthalmoscope (AOSLO) imaging enables visualization of the cones from reflections of waveguiding cone photoreceptors. While there have been significant improvements in automated algorithms for segmenting cones in confocal AOSLO images, the process of labeling data remains labor-intensive and manual. This paper introduces a method based on deep learning (DL) for detecting and segmenting cones in AOSLO images. The models were trained on a semi-automatically labeled dataset of 20 AOSLO batches of images of 18 participants for 0, 1, and 2 from the foveal center. F1 scores were 0.968, 0.958, and 0.954 for 0, 1, and 2, respectively, which is better than previously reported DL approaches. Our method minimizes the need for labeled data by only necessitating a fraction of labeled cones, which is especially beneficial in the field of ophthalmology, where labeled data can often be limited.
@@ 1 Introduction
Considering just the cones, peak foveal density can approach 200,000 cones per $mm^{2}$ [5], making manual labeling impractical. On the other hand, existing automatic labeling techniques may not consistently enable the automatic identification of every cone within an image, particularly in the presence of blood vessels or when the image clarity is compromised. Furthermore, the challenge intensifies when examining retinal locations that are more eccentric from the fovea.
@@ 3.1 Dataset
In this study, we employed a semi-automatically labeled dataset of 20 AOSLO batches of images of 18 healthy participants with normal vision from a wide age range, 14–65 years, representing a wide scope of healthy retinas. Each batch consists of approximately 40 confocal images. The cone centers were first automatically identified with the classical method [7]. Then, missed cones were manually added by a human expert or removed if they were mislabeled by the automatic algorithm, representing approximately 5% of all cones across the dataset images. The data was split on the participant’s level so that images belonging to one participant appeared only in one subset, with a train:test split ratio of 70:30. Therefore, we had 14 batches (540 images) of AOSLO images for training and 6 batches (240 images) for testing. Each image was cropped to $550\times 550$ pixel resolution with 350 labeled cones on each image on average, for a total of 190k segmented cells in the training subset at the starting point. Fig. 1 shows examples of confocal images with labeled cone centers.
Figure 1: Two examples of confocal AOSLO images with labeled cone centers using the existing semi-automatic segmentation method [7] followed by refinement by a medical expert.
@@ 3.2 Human-in-the-loop approach
Fig. 2 illustrates the overall pipeline of the proposed method. AOSLO images were labeled and split into the test and train subsets as described in the dataset section 3.1. To all labeled areas, we applied the Voronoi algorithm to obtain the masks of the cones. Then, the human-in-the-loop step was applied: on the initially labeled AOSLO images, we trained DL-based models to generate semantic masks on unlabeled images. Then, from semantic masks, we calculated the center of mass for each segment (cone), which is basically the center of the cone we manually labeled. Therefore, we could evaluate the models by comparing the obtained centers with ground truth labels. Given the potential for initial inaccuracies, manual correction is a crucial step in the method. This step ensures the precision of the model’s output by allowing the expert to review and adjust the segmented cone centers, mitigating the risk of errors in the initial automated segmentation. Adding new annotations and the manual correction step are not involved in the initial zero iteration, they are only applied starting from the first iteration step.
The Voronoi algorithm is reapplied to the refined data after correction of the centers of the cones, which was done in EXACT [9]. This iterative process refines the segmentation accuracy and enriches the training dataset with additional, corrected instances. Thus, the next 15% of the total number of images is labeled at each iteration, increasing the training dataset. Each iteration concludes with an evaluation step on a test dataset to quantify the improvements. Additionally, at this stage, we apply the $K$-means algorithm for clustering the cones by the mean brightness of the center part (reflection). Therefore, we obtain the percentage of reflecting and non-reflecting cones, which is also a priori information for diagnostics. This cyclical process, encompassing both automated segmentation and expert review, ensures the development of a robust model capable of high-precision cone segmentation.
Figure 2: Pipeline of the method. Voronoi algorithm is applied to initially semi-automatically annotated cones to obtain the masks. Then, a segmentation model was trained, which generates segmentation masks for unlabeled AOSLO images. The center of mass function is applied to get the centers of the cells from segmentation masks. After each iteration step, the model was evaluated using the test subset. A manual correction step is involved in the pipeline to improve the annotations of the segmentation model of initially unlabeled images.
@@ 3.3 Voronoi algorithm
The Voronoi algorithm is one of the more useful geometrical constructions to study point patterns since it provides all the information needed to study proximity relations between points [10]. Connecting surrounding cones and characterizing the number of sides, the Voronoi diagram allows assessment of the degree of hexagonality, and it is often used to show how disease and aging can affect this aspect of packing geometry [1]. In a healthy retina, cones are packed in the most efficient manner possible, which is a hexagonal (honeycomb) arrangement. The degree of hexagonality, therefore, can be used as a proxy for general retinal health. Applying the Voronoi algorithm, we can obtain a reasonably accurate approximation of photoreceptor segmentation by labeling only the centers of cones.
Figure 3: Application of Voronoi algorithm on the labeled AOSLO images: (a) example of the original image; (b) segmented image.
@@ 3.4 Attention-augmented U-Net
Fig. 4 shows the overview of the model. In our model, we applied the concept of flows (vector gradient fields) [20]. This means that we trained a neural network to predict the horizontal and vertical gradients of the topological maps. Additionally, the network predicts a binary map to indicate if a given pixel is inside or outside of regions of interest. Our model was based on the general U-Net architecture [14] with an additional attention-augmented module (AA module) [13]. This module dynamically adjusts the importance of different spatial regions and channels in the input data, enabling the network to prioritize more relevant features for improved segmentation accuracy. Such augmentation facilitates precise localization and detailed segmentation in complex image datasets, which is particularly beneficial in medical imaging applications where accuracy is essential. Attention mechanisms can help the model to focus on relevant features and ignore distractions, therefore, improving segmentation accuracy.
@@
Figure 4: Model overview. (a) Transformation from the center of the cell to a gradient vector field using the Voronoi algorithm. (b) U-Net model with additional Attention-augmented module.
@@ 4 Experiments
The model was trained for 500 epochs on each iteration with stochastic gradient descent with a learning rate of 0.001, a momentum of 0.9, a batch size of 16 images, and a weight decay of 0.0001. All the models were trained on a single NVIDIA A100 graphics processing unit on a machine with two Intel Xeon Gold 6134 3.2 GHz and 96 GB RAM. One training iteration on this setup lasts 30 minutes, with about 10 seconds for the further inference of one image from a batch.
To predict the horizontal and vertical gradients, we used the MSE loss function. We applied the cross-entropy loss function to predict the probability that a pixel was inside or outside a cell.
@@ 4.2 Metrics
where $a$ and $b$ are predicted and ground truth centers, respectively. Detected cones were evaluated using Recall, Precision, and F1-score:
@@ 4.3 Results
Fig. 5 shows an example of the application of the algorithm on the first (a) and second (b) iterations. Green and red circles correspond to ground truth and predicted cell centers, respectively. Yellow squares on the first iteration show unlabeled cells (FNs); on the second iteration, they are correctly labeled.
@@
Figure 5: Evaluation of the proposed method on first (a) and second (b) iterations on a 0 test sample. Green circles correspond to the ground truth cone centers, and red to the predicted centers. Yellow squares show the False Negative predictions of the model.
@@
Table 1 presents the comparative performance of the StarDist, Cellpose, and our models. Recall, Precision, and F1 score were computed separately for 0, 1, and 2 from the fovea. $D_{L_{2}}$ was calculated for all three degrees together. All models are characterized by a fundamental improvement after the second correction: on average, the F1 score improved by 7%. We also see a tendency for the scores to deteriorate slightly at higher eccentricities.
@@
The obtained centers were clustered in terms of brightness to monitor the distribution of light-reflecting and non-reflecting (dark) photoreceptors during training. The $K$-means algorithm is a popular unsupervised machine learning technique for clustering data into a specified number of clusters, denoted by $K$. We applied the $K$-means algorithm with three clusters ($K=3$) during each iteration for additional distribution of reflecting cone control.
@@
Figure 7: Examples of the method performance on the test images located at 0, 1, and 2 from the fovea with applied $K$-Means clustering algorithm. Blue-marked labels correspond to the cones with the highest reflection, green to the lowest, and red to the middle. White boxes show the location of the zoomed area in the right column.
Clustering the cones based on their brightness level is particularly useful in retinal imaging for understanding differences between healthy and diseased retinas. Fig. 7 shows examples of clustering on images of 0, 1, and 2.
@@
Figure 8: The cumulative average number of corrections of cone centers on the first and second iterations per image for Cellpose, StarDist, and the proposed model. The initial training iteration was done with the original labeled dataset; therefore, the number of corrections was equal to zero.
@@ 5 Conclusion
This work describes and evaluates a method for the identification and segmentation of cone photoreceptors from AOSLO confocal images. Models were trained and tested on images covering a more extensive range of images of 18 participants with only 5% labeled cones. Our proposed method received an overall F1 score of 0.968 for cones for 0, 0.958 for 1, and 0.954 for 2, which is better than previously reported DL approaches [4, 3]. Our method can reduce the labeling effort by requiring only a portion of labeled cones and is particularly advantageous in the ophthalmology field, where labeled data can be scarce. The work is limited to the range of degrees of eccentricity from the center from the fovea — 0, 1, and 2. Rods are already present at 1 but peaks in density at around 15; thus, rods become more and more visible on images between cones, which also require detection. Therefore, a potential improvement of the method could be to add the annotations for rods for implementing rods detection for eccentricities more than 2. This could be done using the calculated modality of the AOSLO images.